A halogen gas mixture contained in a 8.0L cylinder holds 25.0g of Cl2 and 10.0g of F2. The temperature of the tank is 25 degree Celsius, what is the total pressure in the cylinder, in atmosphere?
1
Expert's answer
2020-06-22T08:54:04-0400

Moles of "Cl_2=\frac{25}{71}=0.352"

Moles of "F_2=\frac{10}{19}=0.526"

Total moles"(n)=0.526+0.352=0.878"

"T=25\degree C=298K"

Using ideal gas equation,"PV=nRT\implies P\times 8=0.878\times 0.0821\times 298\implies" "P=2.685\ atm"